# Learning Personalized Decision Support Policies

### v2

@@ Appendix E Additional Computational Experiments
$L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$
Synthetic-$2A$ $0.20\pm 0.11$ $0.25\pm 0.08$ $0.15\pm 0.05$ $0.30\pm 0.03$ $0.14\pm 0.05$ $0.31\pm 0.03$ 0.12 $\pm$ 0.04 00 $\pm$ 00
CIFAR-$2A$ $0.25\pm 0.15$ $0.24\pm 0.12$ $0.14\pm 0.04$ $0.34\pm 0.03$ $0.17\pm 0.04$ $0.30\pm 0.04$ $0.14\pm 0.04$ $0.32\pm 0.03$
MMLU-$2A$ $0.18\pm 0.07$ $0.03\pm 0.01$ $0.14\pm 0.04$ $0.03\pm 0.01$ $0.15\pm 0.03$ $0.04\pm 0.01$ $0.15\pm 0.04$ $0.03\pm 0.01$
Table 8: On the CIFAR-$2A$ dataset, we explore the effect of varying the embedding size for KNN, with $K=8$, $W=25$ (Top), and LinUCB (Bottom). We report the expected risk $L_{h}(\pi)$ (lower is better) and expected cost $c(\pi)$ (lower is better) averaged over 5 runs. Recall that in our main paper experiments we used two-dimensional t-SNE embeddings, not these model embeddings, for ease of visualization.
Embedding Size
@@
Table 9: We evaluate $|\mathcal{A}|>3$ in the cost-agnostic setting, $L_{h}(\pi_{t})$ for $t=50,100,250$ for the same setup as Table 1, but with more arms. We find that Modiste works well over time. The success of LinUCB is highly dependent on the embedding space and its geometry.
@@
Figure 11: Snapshots of the learned decision support policies computed at the end of $60$ interactions with randomly sampled users for MMLU-$2A$. The forms of support are colored in t-SNE embedding space. Topics correspond to clusters, clockwise from the top: math, biology, computer science, foreign policy.
@@
In Figure 10, we consider a setting of learning a policy with three forms of support: $\mathcal{A}=\{\textsc{A}_{1},\textsc{A}_{2},\textsc{A}_{3}\}$. Let $c(\textsc{A}_{1})<c(\textsc{A}_{2})=c(\textsc{A}_{3})$. This is an analogous setup to what we do in our human subject experiments in Section Modiste with real users. We indicate the simulated human expertise profiles in the caption. We show how performance on a random sample of 100 points from a held-out set varies with time while learning a decision support policy. In general, we find that the more forms of support to learn, the longer it may take for the online algorithms to “converge.” For various simulations, KNN is more consistent in learning a decision support policy efficiently.
Appendix F Additional User Study Details and Results
CIFAR Algorithm Parameters
For KNN, we take $K=8$, $W=25$, and $\epsilon=0.1$; for both algorithms, we let $\alpha=1$. Cost of either model arm is $0.5$. For each algorithm, we consider two different settings of $\lambda$ ($\lambda\in\{0.75,1.0\}$ and $\in\{0.85,1.0\}$ for KNN and LinUCB respective), selected using the hyper-parameter selection methods described in the main text. [^7] Due to a server-side glitch, 6 of the 125 recruited participants received incorrect feedback on $\leq 2\%$ of trials.
MMLU Algorithm Parameters
We use the same parameters as in the CIFAR task, with the exception that we set the cost of support (i.e., the LLM arm) to $0.1$. Additionally, we employ different $\lambda$$\lambda\in\{0.75,1.0\}$ for KNN and $\lambda\in\{0.95,1.0\}$ for LinUCB – selected using the hyper-parameter selection methods described in the main text.
@@ Appendix F Additional User Study Details and Results
Visualizing Learned Policies. We provide additional details on how the policy generalization snapshots in Figures 4 and 5 are constructed. We save out the parameters learned by each algorithm while a participant interacts with Modiste. After the 60 or 100 trials (depending on which task they participated in), we load in the final state of parameters. We sample the embeddings of unseen points from the respective task dataset (as described in Appendix D) and pass these through the respective algorithm (LinUCB or KNN, depending on which variant the participant was assigned to) – yielding a recommended form of support for said embedding. We color the latent by the form of support. We draw randomly 250 such unseen examples for CIFAR, and 100 for MMLU, when computing the policy recommendations.
We include snapshots for participants, across both tasks, in tha main text. We depict all CIFAR examples and a random sampling of 7 of the 10 participants per variant for MMLU.
Figure 12: Snapshots of the recommended forms of support learned via Modiste for all participants in the CIFAR task, for different $\lambda$. Policies learned using KNN get closer to optimal; see Figure 4.
Cumulative results. We also report in Table 10 the expected loss and cost across all trials. In general, we find similar trends as Table 3.
@@
Table 10: We report expected loss $L_{h}(\pi)$ and expected cost $c(\pi)$ incurred (lower is better) across all trials by Prolific participants for each Algorithm and bold the variant with the lowest $L_{h}(\pi)$. Modiste learns effective, low-cost policies: for CIFAR, Modiste outperforms all baselines. For MMLU, we find that Modiste learns a policy with roughly the same performance as the best offline policy but at half the cost. We also consider different choices of $\lambda$, where $\lambda=1.0$ corresponds to the standard setting and $\lambda\neq 1.0$ corresponds to a cost-aware setting.
Algorithm $L_{h}(\pi)$ $c(\pi)$
H-Only $0.57\pm 0.06$ 00
H-Model $0.44\pm 0.04$ $0.5$
H-Consensus $0.32\pm 0.06$ $0.5$
Population $0.6\pm 0.02$ 00
Modiste (LinUCB, $\lambda=1.0$) $0.28\pm 0.03$ $0.35\pm 0.01$
Modiste (KNN, $\lambda=1.0$) $\mathbf{0.24\pm 0.03}$ $\mathbf{0.39\pm 0.04}$
Modiste (LinUCB, $\lambda=0.85$) $0.37\pm 0.07$ $0.25\pm 0.01$
Modiste (KNN, $\lambda=0.75$) $0.24\pm 0.03$ $0.4\pm 0.03$
Algorithm $L_{h}(\pi)$ $c(\pi)$
H-Only $0.44\pm 0.07$ 00
H-LLM $\mathbf{0.22\pm 0.06}$ $\mathbf{0.1}$
Population $0.34\pm 0.08$ $0.05\pm 0.00$
Modiste (LinUCB, $\lambda=1.0$) $0.26\pm 0.07$ $0.05\pm 0.00$
Modiste (KNN, $\lambda=1.0$) $0.32\pm 0.1$ $0.06\pm 0.01$
Modiste (LinUCB, $\lambda=0.95$) $0.31\pm 0.09$ $0.06\pm 0.01$
Modiste (KNN, $\lambda=0.75$) $0.32\pm 0.1$ $0.06\pm 0.01$

### v3

@@ Appendix D Additional Computational Experiments
Synthetic-$2A$ $0.20\pm 0.11$ $0.15\pm 0.05$ $0.14\pm 0.05$ 0.12 $\pm$ 0.04
CIFAR-$2A$ $0.25\pm 0.15$ $0.14\pm 0.04$ $0.17\pm 0.04$ $0.14\pm 0.04$
MMLU-$2A$ $0.18\pm 0.07$ $0.14\pm 0.04$ $0.15\pm 0.03$ $0.15\pm 0.04$
Table 4: On the CIFAR-$2A$ dataset, we explore the effect of varying the embedding size for KNN, with $K=8$, $W=25$ (Top), and LinUCB (Bottom). We report the expected risk $L_{h}(\pi)$ (lower is better) averaged over 5 runs. Recall that in our main paper experiments we used two-dimensional t-SNE embeddings, not these model embeddings, for ease of visualization.
Embedding
Size
@@
Table 5: We evaluate $|\mathcal{A}|>3$ via $L_{h}(\pi_{t})$ for $t=50,100,250$ for the same setup as Table 1, but with more arms. We find that Modiste works well over time. The success of LinUCB is highly dependent on the embedding space and its geometry.
@@
Figure 10: Snapshots of the learned decision support policies computed at the end of $60$ interactions with randomly sampled users for MMLU-$2A$ with LinUCB policies displayed atop KNN ones. The forms of support are colored in t-SNE embedding space. Topics correspond to clusters, clockwise from the top: math, biology, computer science, foreign policy.
@@
In Figure 9, we consider a setting of learning a policy with three forms of support: $\mathcal{A}=\{\textsc{A}_{1},\textsc{A}_{2},\textsc{A}_{3}\}$. This is an analogous setup to what we do in our human subject experiments in Section 7. We indicate the simulated human expertise profiles in the caption. We show how performance on a random sample of 100 points from a held-out set varies with time while learning a decision support policy. In general, we find that the more forms of support to learn, the longer it may take for the online algorithms to “converge.” For various simulations, KNN is more consistent in learning a decision support policy efficiently.
Appendix E Additional User Study Details and Results
Figure 12: Snapshots of the recommended forms of support learned via Modiste for all participants in the CIFAR task. Policies learned using KNN get closer to optimal; see Figure 8.
Algorithm Parameters
For KNN, we take $K=8$, $W=25$, and $\epsilon=0.1$; for both algorithms, we let $\alpha=1$.  [^6] Due to a server-side glitch, 6 of the CIFAR 125 recruited participants received incorrect feedback on $\leq 2\%$ of trials. We use the same parameters for both tasks.
@@ Appendix E Additional User Study Details and Results
Visualizing Learned Policies. We provide additional details on how the policy generalization snapshots in Figures 8 and 5 are constructed. We save out the parameters learned by each algorithm while a participant interacts with Modiste. After the 60 or 100 trials (depending on which task they participated in), we load in the final state of parameters. We sample the embeddings of unseen points from the respective task dataset (as described in Appendix C) and pass these through the respective algorithm (LinUCB or KNN, depending on which variant the participant was assigned to) – yielding a recommended form of support for said embedding. We color the latent by the form of support. We draw randomly 250 such unseen examples for CIFAR, and 100 for MMLU, when computing the policy recommendations.
We include snapshots for participants, across both tasks, in the main text. We depict all CIFAR examples and a random sampling of 7 of the 10 participants per variant for MMLU.
@@
Appendix F Further Discussion
Our human subject experiments validate trends observed in the computational experiments, demonstrating that Modiste can be used to learn decision support policies online. We find that learning personalized policies leads to performance benefits when encountering users with “varying” expertise profiles, and recovers fixed policies for participants with “strictly better” profiles. Future work modeling human idiosyncrasies (Steyvers and Kumar 2022) explicitly in synthetic decision-makers may bolster the effectiveness of personalization.